The length of a rectangle is 5 meters less than 4 times the width. If the perimeter is 150 meters, find the length and the width of the rectangle.
Answer by tiffany222(56)

Let width be x
Let length be y
y = 4x-5
P = 2(width) + 2(length) or P=2x + 2y
P = 2x + 2y
150 = 2x + 2(4x-5)
150 = 2x + 8x - 10
150 = 10x - 10
160 = 10x
16 = x
y = 4x-5 = 4(16) - 5 = 59
Final Answer:
Width = 16 m
Length = 59 m